# 3D Shape Completion on Unseen Categories: A Weakly-supervised Approach

### v1

@@ III-B Category-specific Shape Refinement
According to the idea of self-supervised point cloud completion methods [23, 24, 25], complete shapes can be inferred from the partial shapes of the target categories. These methods take a partial input and use partial shapes of the input object along with the partial matching loss to guide the model in reconstructing the complete shape. In this way, the model can discover many common patterns belonging to the shapes of the category and thus infer the complete shapes of the input. As shown in Eq. (3), the point-based partial matching loss is a unidirectional Chamfer Distance loss, which guides the existing parts of the predicted shape to be close to the input partial shape.
@@
where $\mathbf{P}_{in}$ and $\mathbf{P}$ denote the partial point cloud and its corresponding ground truth, respectively.
Unfortunately, directly applying this loss to voxels is impracticable. Unlike point-cloud data, voxel representations do not impose constraints on the number of foreground voxels. Directly applying the partial L1 loss to voxels tends to cause most of the voxels to be predicted as the foreground, as shown in Figure 7.
To solve this problem, we propose a voxel-based partial matching loss. First, we compute the partial L1 loss to constrain the existing parts of the predicted shape to be close to the input partial shape. Denote $\mathbf{S}$ the partial scan and $\mathbf{O}$ the refined result. The resolution of $\mathbf{S}$ and $\mathbf{O}$ is $N$, and the voxel value is 1 if this voxel is occupied otherwise the voxel value is 0. The partial L1 loss $L_{p}$ is computed as shown in Formula 4.
@@
Although the partial L1 loss can constrain the existing parts of the predicted result to be close to the partial input, it cannot guide the result to be a reasonable complete shape. Due to the lack of constraints on the number of occupied voxels, the model tends to fill most voxels with “1”. The point-based partial matching loss can successfully guide the predicted shape because the predicted point number is fixed. That means the point cloud completion models should only predict the positions of a fixed number of points. Inspired by this, we fix the number of foreground voxels for all predicted results. Specifically, we design an occupancy loss that sums up all the predicted voxel values and make them close to a predefined occupied voxel number $H$. Since the number of occupied voxels varies across different objects, a fixed value of $H$ for all objects may significantly compromise shape integrity or result in redundancy, particularly if $H$ deviates significantly from the actual number of occupied voxels. To address this issue, we adaptively set $H$ of the complete target shape to be $\alpha$ times the occupied voxel number of the corresponding partial input.
$\begin{split}L_{s}=|\sum_{i=0}^{N^{3}}\mathbf{O}_{i}-H|\,,\\ H=\alpha\sum_{i=0}^{N^{3}}\mathbf{S}_{i}\,.\end{split}$ (5)
Additionally, $L_{s}$ can calculate the actual number of occupied voxels only when the voxel value is 0 or 1. However, in practice, the value of each voxel outputted by the network falls between 0 and 1 due to the Softmax function. To make the predicted voxel values close to 0 or 1, we propose a variance loss that computes the variance of the predicted voxel values and make them larger as much as possible.
@@
The voxel-based partial matching loss is the weighted sum of partial L1 loss, occupancy loss and variance loss:
@@
While the voxel-based partial matching loss facilitates the model to reconstruct the missing parts of the inputs by learning the common patterns of each category, the model may predict significantly deviated results in these regions due to the lack of supervision. On the other hand, the coarse shapes, while not completely accurate, do exhibit a general similarity to the complete shapes of the inputs. In order to guide the model to predict more reasonable missing parts for the partial input, we use the missing parts of the coarse shapes outputted by PSLN as weak supervision. Specifically, we first remove the regions of the coarse shape that overlap with the partial input. We refer to the remaining part as the missing part and denote it as $\mathbf{T}$. Then we compute the partial L1 loss for $\mathbf{T}$ and the predicted result. We term this loss as coarse shape loss and denote it as $L_{m}$. Note that we only assign a small weight $\lambda$ for this loss term. Overall, the total loss of our CSRN is:
@@
We adopt the same framework as PSLN to refine our coarse outputs. The inputs are coarse shapes outputted by PSLN and category-specific priors. PatchComplete only utilizes fixed priors from seen categories to infer the shapes of unseen categories. However, these fixed priors are not easily adaptable to all unseen data categories since each data category has its unique shape characteristics. For example, lamps typically exhibit a stick-shaped lampstand, while desks commonly feature a flat desktop and pillar-shaped legs. To address this issue, we introduce category-specific priors in the refinement stage, replacing the fixed priors derived from the training data. Specifically, to make use of more partial shapes, we first calculate the overlap area between each pair of partial shapes from the target category. Then we select M pairs with the least overlap areas and concatenate them to construct the prior bank. Note that all partial shapes within this bank originate from distinct objects. In this way, the prior bank for each category contains various structures that are specific to that category. We visualize some priors in Figure 2.
@@ IV-A1 Dataset
To ensure a comprehensive evaluation of our method, we have conducted experiments on both synthetic and real-world data. The synthetic data we utilize are derived from ShapeNet [53]. The partial shapes are obtained by scanning the objects from ShapeNet at random angles. Following PatchComplete, we train our model with data from 18 categories (i.e., table, chair, sofa, cabinet, clock, bookshelf, piano, microwave, stove, file cabinet, trash bin, bowl, display, keyboard, dishwasher, washing machine, pots, faucet, and guitar) and subsequently test it on data from other 8 novel categories (i.e., bathtub, lamp, bed, bag, printer, laptop, bench, and basket). The training set and testing set comprises 3202 and 1325 models, respectively, each of which consisted of 4 partial scans.
For real-world data, we used real scans from ScanNet [54], which provides both partial scans and their complete ground truth. In this dataset, the training set contains 7537 models from 8 categories (i.e., chair, table, sofa, trash bin, cabinet, bookshelf, file cabinet, and monitor), while the testing set contains 1191 models from the remaining 6 categories (i.e., bathtub, lamp, bed, bag, basket, and printer). Each model has only one partial scan.
The partial and complete shapes from both synthetic and real-world data are represented as voxel grids with the resolution of $32^{3}$.
@@ IV-A2 Evaluation Metric
We use the Intersection over Union (IoU) and F1 score to evaluate the similarity between the predicted voxel grids and the ground truth. The IoU measures the overlap between two sets of data and the F1 score is the harmonic mean of precision and recall. In addition, following [26], we convert the predicted voxel grids to point clouds and then employ the Chamfer distance (CD) to compare the differences between the predicted point sets and the ground truth.
@@ IV-B Implementation Details
Our model is trained from scratch on both the ShapeNet and ScanNet datasets, without leveraging any pre-trained models. Four partial scans are utilized to compute the voxel-based partial matching loss for ShapeNet, while one partial scan is used for ScanNet. We used Adam Optimizer to train our model. The batch size is 10. The initial learning rate of PSLN and CSRN is 1e-3 and 1e-4, respectively. The learning rate exponentially decays by 0.5 for every 50 epochs and the models are trained for 120 epochs for both stages. The training process is conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
We train the PSLN on the unseen categories and determine the hyperparameters based on the IoU between the predicted results and the corresponding complete shapes from the seen categories. Subsequently, we train the CSRN on the unseen categories and select the hyperparameters using the IoU between the predicted results and the partial inputs, as complete shapes are typically unavailable for unseen categories in practical scenarios. The complete shapes from the unseen categories of our dataset are solely employed for evaluation purposes. In Section IV-C5, we present experimental results to validate the efficacy of this strategy.
@@ IV-C Comparison with State-of-the-Art Methods
Figure 3: Comparison of the visual results on ShapeNet dataset. From top to bottom, the categories of the objects are bag, lamp, bathtub, bed, basket, printer, laptop and bench. (a) Partial input (b) AutoSDF. (c) IFNet. (d) Fewshot. (e) PatchComplete. (f) Ours. (g) Ground truth.
Figure 4: Comparison of the visual results on ScanNet dataset. From top to bottom, the categories of the objects are bag, lamp, bathtub, bed, basket and printer. (a) Partial input (b) AutoSDF. (c) IFNet. (d) Fewshot. (e) PatchComplete. (f) Ours. (g) Ground truth.
@@ IV-C1 Results on synthetic data
We first conduct experiments on ShapeNet. Table I shows the quantization results of our method and other state-of-the-art (SOTA) shape completion and reconstruction methods, including Few-Shot [30], IF-Nets [29], AutoSDF [8] and PatchComplete [9]. We can see that our method achieves the best results among all categories in terms of IoU and F1 score. Specifically, the average IoU and F1 of our model are almost double those of Few-Shot, IF-Nets and AutoSDF. Compared to PatchComplete, our model improves the average IoU and F1 score by 0.045 and 0.043, respectively. Especially, the IoU of our model surpasses that of PatchComplete by more than 0.05 in four categories, i.e., bag, bed, printer and laptop. In terms of the CD metric, our method achieves the best average result and exceeds the second-place method by 0.33. Besides, our method outperforms all other methods in six categories and only slightly falls behind PatchComplete in the remaining two categories.
Figure 3 displays the qualitative results of our method and the baselines on the ShapeNet dataset. Note that the objects depicted in these images are meshes converted from predicted voxels and have undergone smoothing. As shown in Figure 3, our method is able to recover the missing part effectively while preserving the partial input. AutoSDF and Fewshot tend to modify the partial and face challenges in accurately predicting missing parts. IFNet is susceptible to generating noises during reconstruction that do not belong to the objects themselves. And the shape reconstructed by PatchComplete is incomplete, with some holes resulting from missing details. However, our method can reconstruct a more complete shape and restore finer details, such as the corners of the printer and the basket.
@@ IV-C2 Results on real-world data
Our method is effective not only on synthetic datasets but also demonstrates strong performance on real-world datasets. Table II shows the quantization results of our method and other SOTA shape completion and reconstruction methods. Overall, our method outperforms all baselines by large margins in terms of average IoU, F1 score and CD, especially with a nearly 16% lead over the second-ranked method in terms of IoU metric. In addition, our method yields a comprehensive lead over other methods in all categories in terms of CD, while the IoU and F1 score are only slightly behind those of PatchComplete in one category. It is worth noting that our method’s IoU and F1 score in the bathtub and bed categories both demonstrate a lead of more than 0.1 over those of the second-ranked method.
Figure 4 depicts the visual results of our method and the baselines on ScanNet dataset. The visual results indicate that all methods perform similarly on the ScanNet dataset as they did on the ShapeNet dataset. Our method is capable of predicting missing parts that are more reasonable and finely detailed, while more effectively preserving the input parts compared to other methods.
@@ IV-C3 Comparison with point cloud completion methods
To provide a more comprehensive evaluation of our method, we also compared it with three SOTA point cloud completion approaches, i.e., PCN [32], GRNet [36] and AnchorFormer [55]. For a fair comparison, the number of output points is set to 2048, which is close to the mean point number (1922) of the complete point clouds converted from voxel grids in our dataset. All methods are trained using data from seen categories and subsequently tested on unseen categories. As depicted in Table III, our method outperforms all other point cloud completion methods across all categories on ShapeNet data. In terms of average CD, our method surpasses the second-ranked method by 1.62. Regarding the ScanNet dataset, as shown in Table IV, our method significantly outperforms point cloud completion methods in 5/6 categories on ScanNet dataset, and the average CD is 0.4 lower than the AnchorFormer. Figure 5 presents the visual results of our model and point cloud approaches. It is evident that our method accurately reconstructs the complete shapes while recovering the details, such as the corners of the bathtub and the legs of the bed. Moreover, our model generates smoother surfaces than the compared point cloud methods.
Figure 5: Comparison of visual results between our methods and point cloud completion methods. (a) Partial input, (b) GRNet [36], (c) AnchorFormer [55], (d) Ours, (e) Ground truth
@@ IV-C4 Analysis of computational complexity and model size
Table V shows the FLOPs (Floating Point Operations) and model sizes of our PSLN and PatchComplete. It can be seen that our PSLN achieves better results than PatchComplete while requiring only half the number of FLOPs and one-fourth of the model size. Moreover, our PSLN learns the correlation between the priors and input at the feature level and can be trained end-to-end, whereas PatchComplete requires a four-time training.
TABLE V: Comparison of FLOPs and model size. The best results are highlighted in bold.
@@
PatchComplete 1335.1 697.7 0.429
PSLN 652.1 180.8 0.441
Figure 6: The IoU curve during the training process of CSRN on ShapeNet dataset. The y-axis represents the IoU between the predicted complete shape and the partial input, while the x-axis represents the training epoch. (a) bag, (b) lamp, (c) bathtub, (d) bed, (e) basket, (f) printer, (g) laptop, (h) bench.
@@
PSLN w/o MSL 0.409 0.380 0.355 0.450 0.345 0.453 0.368 0.424 0.497 0.331 0.322 0.347 0.361 0.319 0.352 0.286
PSLN w/ MSL 0.441 0.391 0.380 0.525 0.383 0.493 0.392 0.472 0.492 0.385 0.354 0.362 0.438 0.369 0.445 0.343
PSLN+CSRN 0.474 0.450 0.380 0.542 0.426 0.509 0.458 0.533 0.492 0.432 0.384 0.362 0.529 0.448 0.523 0.343

### v2

@@ III-B Category-specific Shape Refinement
According to the idea of self-supervised point cloud completion methods [23, 24, 25], complete shapes can be inferred from the partial shapes of the target categories. These methods take a partial input and use partial shapes of the input object along with the partial matching loss to guide the model in reconstructing the complete shape. In this way, the model can discover many common patterns belonging to the shapes of the category and thus infer the complete shapes of the input. Intuitively, one can apply the point-based partial matching loss, which is a unidirectional CD loss defined as
@@
to promote the existing parts of the predicted shape $\mathbf{P}$ to be close to the input partial shape $\mathbf{P}_{in}$. However, unlike the point-based representation, voxel representations do not impose constraints on the number of occupied voxels, and directly applying the partial L1 loss to voxel data tends to result in most voxels being occupied (see Figure 7).
To this end, we propose a voxel-based partial matching loss. First, we compute the partial L1 loss to constrain the existing parts of the predicted shape to be close to the input partial shape. Denote $\mathbf{S}$ the partial scan and $\mathbf{O}$ the refined result. The resolution of $\mathbf{S}$ and $\mathbf{O}$ is $N$, and the voxel value is 1 if this voxel is occupied; otherwise, the voxel value is 0. The partial L1 loss $L_{p}$ is computed as
@@
Although the partial L1 loss can constrain the existing parts of the predicted result to be close to the partial input, it cannot guide the result to be a reasonable and complete shape. Due to the lack of constraints on the number of occupied voxels, the model tends to fill most voxels with “1”. The point-based partial matching loss can successfully guide the predicted shape because the predicted point number is fixed. That means the point cloud completion models should only predict the positions of a fixed number of points. Inspired by this, we fix the number of foreground voxels for all predicted results. Specifically, we design an occupancy loss that sums up all the predicted voxel values and makes them close to a predefined occupied voxel number $H$. Since the number of occupied voxels varies across different objects, a fixed value of $H$ for all objects may significantly compromise shape integrity or result in redundancy, particularly if $H$ deviates significantly from the actual number of occupied voxels. To address this issue, we adaptively set $H$ of the complete target shape to be $\alpha$ times the occupied voxel number of the corresponding partial input:
$L_{s}=|\sum_{i=0}^{N^{3}}\mathbf{O}_{i}-H|,{\rm with}~{}H=\alpha\sum_{i=0}^{N^{3}}\mathbf{S}_{i}.$ (5)
Additionally, $L_{s}$ can calculate the actual number of occupied voxels only when the voxel value is 0 or 1. However, in practice, the value of each voxel outputted by the network falls between 0 and 1 due to the Softmax function. To make the predicted voxel values close to 0 or 1, we propose a variance loss that computes the variance of the predicted voxel values and makes them larger as much as possible:
@@
The voxel-based partial matching loss is the weighted sum of partial L1 loss, occupancy loss, and variance loss:
@@
While the voxel-based partial matching loss facilitates the model in reconstructing the missing parts of the inputs by learning the typical patterns of each category, the model may predict significantly deviated results in these regions due to the lack of supervision. On the other hand, the coarse shapes, while not completely accurate, do exhibit a general similarity to the complete shapes of the inputs. In order to guide the model in predicting more reasonable missing parts for the partial input, we use the missing parts of the coarse shapes outputted by CoSL as weak supervision. Specifically, we first remove the regions of the coarse shape that overlap with the partial input. We refer to the remaining part as the missing part and denote it as $\mathbf{T}$. Then, we compute the partial L1 loss for $\mathbf{T}$ and the predicted result. We term this loss as coarse shape loss and denote it as $L_{m}$. Note that we only assign a small weight $\lambda$ for this loss term. Overall, the total loss of our CaSR is:
@@
We adopt the same framework as PSLN in the CoSL stage to refine our coarse outputs. The inputs are coarse shapes outputted by CoSL and category-specific priors. PatchComplete only utilizes fixed priors from seen categories to infer the shapes of unseen categories. However, these fixed priors are not easily adaptable to all unseen data categories since each data category has its unique shape characteristics. For example, lamps typically exhibit a stick-shaped lampstand, while desks commonly feature a flat desktop and pillar-shaped legs. To address this issue, we introduce category-specific priors in the refinement stage, replacing the fixed priors derived from the training data. Specifically, to make use of more partial shapes, we first calculate the overlapping area between each pair of partial shapes from the target category. Then, we select M pairs with the least overlapping areas and concatenate them to construct the prior bank. Note that all partial shapes within this bank originate from distinct objects. In this way, the prior bank for each category contains various structures that are specific to that category. We visualize some priors in Figure 2.
@@ IV-A1 Dataset
To ensure a comprehensive evaluation of our method, we have conducted experiments on both synthetic and real-world data. The synthetic data we utilize are derived from ShapeNet [55]. The partial shapes are obtained by scanning the objects from ShapeNet at random angles. Following PatchComplete, we train our model with data from 18 categories (i.e., table, chair, sofa, cabinet, bookshelf, piano, microwave, stove, file cabinet, trash bin, bowl, display, keyboard, dishwasher, washing machine, pots, faucet, and guitar) and subsequently test it on data from other 8 unseen categories (i.e., bathtub, lamp, bed, bag, printer, laptop, bench, and basket). The training set and test set comprise 3202 and 1325 models, respectively, each of which consisted of 4 partial shapes scanned from random viewpoints.
For real-world data, we used real scans from ScanNet [56], which provides both partial scans and their complete ground truth. In this dataset, the training set contains 7537 models from 8 categories (i.e., chair, table, sofa, trash bin, cabinet, bookshelf, file cabinet, and monitor), while the test set contains 1191 models from 6 unseen categories (i.e., bathtub, lamp, bed, bag, basket, and printer). Each model has only one partial scan.
The partial and complete shapes from both synthetic and real-world data are represented as voxel grids with the resolution of $32\times 32\times 32$.
@@ IV-A2 Evaluation Metric
We use the Intersection over Union (IoU) and F1 score to evaluate the similarity between the predicted voxel grids and the ground truth. The IoU measures the overlap between two sets of data, and the F1 score is the harmonic mean of precision and recall. In addition, following [26], we convert the predicted voxel grids to point clouds and then employ the Chamfer distance (CD) to compare the differences between the predicted point sets and the ground truth.
@@ IV-B Implementation Details
Our model is trained from scratch on both the ShapeNet and ScanNet datasets without leveraging any pre-trained models. Four partial scans are utilized to compute the voxel-based partial matching loss for ShapeNet, while one partial scan is used for ScanNet. We used Adam Optimizer to train our model. The batch size is 10. The initial learning rate of CoSL and CaSR is 1e-3 and 1e-4, respectively. The learning rate exponentially decays by 0.5 for every 50 epochs, and the models are trained for 120 epochs for both stages. The training process is conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
We train the CoSL on the unseen categories and determine the hyperparameters based on the IoU between the predicted results and the corresponding complete shapes from the seen categories. Subsequently, we train the CaSR on the unseen categories and select the hyperparameters using the IoU between the predicted results and the partial inputs, as complete shapes are typically unavailable for unseen categories in practical scenarios. The complete shapes from the unseen categories of our dataset are solely employed for evaluation purposes. In Section IV-C5, we present experimental results to validate the efficacy of this strategy.
@@ IV-C Comparison with State-of-the-Art Methods
Figure 3: Comparison of the visual results on the ShapeNet dataset. From top to bottom, the categories of the objects are bag, lamp, bathtub, bed, basket, printer, laptop, and bench. (a) Partial input (b) AutoSDF. (c) IFNet. (d) Fewshot. (e) SDFusion. (f) PatchComplete. (g) Ours. (h) Ground truth.
Figure 4: Comparison of the visual results on ScanNet dataset. From top to bottom, the categories of the objects are bag, lamp, bathtub, bed, basket, and printer. (a) Partial input (b) AutoSDF. (c) IFNet. (d) Fewshot. (e) SDFusion. (f) PatchComplete. (g) Ours. (h) Ground truth.
@@ IV-C1 Results on synthetic data
We first conducted experiments on ShapeNet. Table I shows the quantization results of our method and other state-of-the-art (SOTA) shape completion and reconstruction methods, including Few-Shot [30], IF-Nets [29], AutoSDF [8], SDFusion [31] and PatchComplete [9]. Both our method and PatchComplete take as input a partial shape along with a set of priors, in contrast to other approaches that only use a single partial shape. However, the priors are selected from the training set without requiring additional data. We can see that our method achieves the best results among all categories in terms of IoU and F1 score. Specifically, the average IoU and F1 of our model are almost double those of Few-Shot, IF-Nets, and AutoSDF. Compared to PatchComplete, our model improves the average IoU and F1 score by 0.045 and 0.043, respectively. Especially the IoU of our model surpasses that of PatchComplete by more than 0.05 in four categories, i.e., bag, bed, printer, and laptop. In terms of the CD metric, our method achieves the best average result and exceeds the second-place method by 0.33. Besides, our method outperforms all other methods in six categories and only slightly falls behind PatchComplete in the remaining two categories.
Figure 3 displays the qualitative results of our method and the baselines on the ShapeNet dataset. Note that the objects depicted in these images are meshes converted from predicted voxels and have undergone smoothing. As shown in Figure 3, our method is able to recover the missing part effectively while preserving the partial input. AutoSDF and Few-Shot tend to modify the partial and face challenges in accurately predicting missing parts. Although SDFusion better preserves the partial input than AutoSDF and Few-Shot, it struggles to recover missing regions. IFNet is susceptible to generating noises during reconstruction that do not belong to the objects themselves. The shape reconstructed by PatchComplete is incomplete, with some holes resulting from missing details. However, our method can reconstruct a more complete shape and restore finer details, such as the corners of the printer and the basket.
@@ IV-C2 Results on real-world data
Our method is effective not only on synthetic datasets but also demonstrates strong performance on real-world datasets. Table II shows the quantization results of our method and other SOTA shape completion and reconstruction methods. Overall, our method outperforms all baselines by large margins in terms of average IoU, F1 score, and CD, especially with a nearly 16% lead over the second-ranked method in terms of IoU metric. In addition, our method yields a comprehensive lead over other methods in all categories in terms of CD, while the IoU and F1 score are only slightly behind those of PatchComplete in one category. It is worth noting that our method’s IoU and F1 score in the bathtub and bed categories both demonstrate a lead of more than 0.1 over those of the second-ranked method.
Figure 4 depicts the visual results of our method and the baselines on the ScanNet dataset. The visual results indicate that all methods perform similarly on the ScanNet dataset as they did on the ShapeNet dataset. Our method is capable of predicting missing parts that are more reasonable and finely detailed while more effectively preserving the input parts compared to other methods.
@@ IV-C3 Comparison with point cloud completion methods
To provide a more comprehensive evaluation of our method, we also compared it with three SOTA point cloud completion approaches, i.e., PCN [32], GRNet [36] and AnchorFormer [57]. For a fair comparison, the number of output points is set to 2048, which is close to the mean point number (1922) of the complete point clouds converted from voxel grids in our dataset. All methods were trained using data from seen categories and subsequently tested on unseen categories. As depicted in Table III, our method outperforms all other point cloud completion methods across all categories on ShapeNet data. In terms of average CD, our method surpasses the second-ranked method by 1.62. Regarding the ScanNet dataset, as shown in Table IV, our method significantly outperforms point cloud completion methods in 5/6 categories on the ScanNet dataset, and the average CD is 0.4 lower than the AnchorFormer. Figure 5 presents the visual results of our model and point cloud approaches. It is evident that our method accurately reconstructs the complete shapes while recovering the details, such as the corners of the bathtub and the legs of the bed. Moreover, our model generates smoother surfaces than the compared point cloud methods.
Figure 5: Comparison of visual results between our methods and point cloud completion methods. (a) Partial input, (b) GRNet [36], (c) AnchorFormer [57], (d) Ours, (e) Ground truth.
@@ IV-C4 Analysis of computational complexity and model size
Table V shows the FLOPs (Floating Point Operations) and model sizes of our PSLN and PatchComplete under different resolutions of the used data. It is evident that at a resolution of 32, our PSLN achieves superior results compared to PatchComplete while consuming only half the number of FLOPs and a quarter of the model size. Increasing to a resolution of 64, the FLOPs of our PSLN decrease to approximately 1/25 of those of PatchComplete, accompanied by a model size that is merely 1/7. The increase in model size is only 6M in comparison to the values observed at a resolution of 32. These minimal increases are attributed to our model’s ability to downsample priors and partial inputs at the feature level, reducing the computational cost of attention maps. Moreover, our PSLN can be trained end-to-end, whereas PatchComplete requires four-time training.
TABLE V: Comparison of FLOPs and model size.“Res.” means the resolution of the used voxel grids. The best results are highlighted in bold.
@@
PatchComplete (Res. 32) 1335.1 697.7 0.429
PSLN (Res. 32) 652.1 180.8 0.441
PatchComplete (Res. 64) 30748.5 1339.1 0.562
PSLN (Res. 64) 1235.8 186.7 0.598
Figure 6: The IoU curve during the training process of CaSR on the ShapeNet dataset. The y-axis represents the IoU between the predicted complete shape and the partial input, while the x-axis represents the training epoch. (a) bag, (b) lamp, (c) bathtub, (d) bed, (e) basket, (f) printer, (g) laptop, (h) bench.
@@
CoSL w/o MSL 0.409 0.380 0.355 0.450 0.345 0.453 0.368 0.424 0.497 0.331 0.322 0.347 0.361 0.319 0.352 0.286
CoSL w/ MSL 0.441 0.391 0.380 0.525 0.383 0.493 0.392 0.472 0.492 0.385 0.354 0.362 0.438 0.369 0.445 0.343
CoSL+CaSR 0.474 0.450 0.380 0.542 0.426 0.509 0.458 0.533 0.492 0.432 0.384 0.362 0.529 0.448 0.523 0.343